# DiCoTTA: Domain-invariant Learning for Continual Test-time Adaptation

### v1

@@ 3.3 Domain-invariant learning
DiCoTTA updates the model through domain-invariant learning and self-training. For domain-invariant learning, we employ a two-stage process. In the first stage, the domain information extractor and amplifier are updated to extract domain-specific information, i.e., domain embeddings, from the encoder features. In the second stage, the encoder is updated to be domain-invariant by aligning the current domain embeddings with the previous domain prototypes (Sec. 3.2). Alternating these two stages allows the encoder to gradually become domain-invariant, and the domain information extractor and amplifier to become sensitive to the remaining domain-specific information of the encoder. In addition, following previous work [71, 17, 44], we update our model by a self-training loss $\mathcal{L}_{\textrm{self}}$, a cross-entropy loss between our model’s prediction $\hat{y}$ and a pseudo label $\tilde{y}$ generated by its exponential moving average [66]:
@@
where $C$ is the total number of classes. For inference, we use $\hat{y}$ as the classification prediction. Details of the two stages are described below.
Figure 3: Empirical analysis on the impact of domain-invariant learning in DiCoTTA. (a) 2D visualization of the distribution of domain embeddings from different domains before and after DiCoTTA. (b) The domain gap between different domains measured by Chamfer distance before and after applying DiCoTTA.
Figure 4: Empirical analysis on domain embeddings and prototypes. (a) Analysis of class classification and domain classification accuracy using domain embeddings during CTTA iterations. (b) 2D visualization of domain embeddings and domain prototypes. (c) MMD between the target and original prototypes before the update and that between the target and updated prototypes.
Stage 1: Extracting domain embeddings. To maintain the semantic information of the encoder features during domain-invariant learning, we first extract their domain-specific information in the form of embedding vectors. To this end, we introduce the domain information extractor that extracts domain embeddings from output features of the encoder, and the domain amplifier that amplifies the domain-specific information of the encoder to facilitate the extractor. Let $\mathcal{F}_{\textrm{cur}}=\{f_{i}\}_{i=1}^{n}$ be the domain embeddings of test input images from the current test domain and $\mathcal{P}_{\textrm{pre}}$ the previous domain’s prototypes. The domain information extractor and amplifier learn a domain embedding space by distinguishing $\mathcal{F}_{\textrm{cur}}$ and $\mathcal{P}_{\textrm{pre}}$ with a domain discrimination loss:
$\mathcal{L}_{\textrm{dis}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})=$
$-\frac{1}{n}\bigg{\{}\displaystyle\sum_{f_{i}\in\mathcal{F}_{\textrm{cur}},}$ $\,\log(D(f_{i}))-\sum_{p_{i}\in\mathcal{P}_{\textrm{pre}}}\log(1-D(p_{i}))\bigg{\}},$ (5)
@@
Stage 2: Learning domain-invariant features. The second stage aims to optimize the encoder to achieve domain invariance across diverse test domains alongside classification (Eq. 4). DiCoTTA accomplishes this by aligning the current domain embeddings and the previous domain’s prototypes. To be specific, we update the encoder and its classification head while freezing the domain information extractor and amplifier. Given current domain embeddings $\mathcal{F}_{\textrm{cur}}$ and previous domain prototypes $\mathcal{P}_{\textrm{pre}}$, the domain-invariant learning loss is given by
@@
By minimizing this loss, the encoder’s features are trained to produce domain embeddings that are indistinguishable between different domains, making them domain-agnostic and generalizable to future unseen domains.
@@ 3.4 Domain prototype update
Since the model used to extract the domain prototypes is constantly updated during CTTA, the domain prototypes extracted from the earlier model become outdated. Thus, the domain prototypes should also be updated to reflect the update of the model, approximating prototypes extracted from the updated model. Inspired by Asadi et al. [2], we aim to preserve the set similarity between the previous domain prototypes $\mathcal{P}_{\textrm{pre}}$ and the current domain embeddings $\mathcal{F}_{\textrm{cur}}$ to effectively keep the relationship between them intact. In specific, we measure the Chamfer distance between the two sets $\mathcal{P}_{\textrm{pre}}$ and $\mathcal{F}_{\textrm{cur}}$. For iteration $t$, we update the domain prototypes $\mathcal{P}_{\textrm{pre}}^{t-1}$ to $\mathcal{P}_{\textrm{pre}}^{t}$ to ensure that the Chamfer distance between previous domain prototypes and current domain embeddings remains consistent before and after the model update. Specifically, the domain prototypes $\mathcal{P}_{\textrm{pre}}^{t}$ are updated by minimizing the prototype updating loss:
@@
$d_{\textrm{CD}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})$ $\,=\sum_{f_{i}\in\mathcal{F}_{\textrm{cur}}}\min_{p_{j}\in\mathcal{P}_{\textrm{pre}}}\left\|p_{j}-f_{i}\right\|_{2}^{2}$
$\,+\sum_{p_{j}\in\mathcal{P}_{\textrm{pre}}}\min_{f_{i}\in\mathcal{F}_{\textrm{cur}}}\left\|p_{j}-f_{i}\right\|_{2}^{2}.$ (8)
@@
3.5 Empirical justification
Domain-invariant learning. We analyze the impact of domain-invariant learning both qualitatively and quantitatively. Fig. 3(a) provides a $t$-SNE visualization [67] of domain embeddings from different domains (i.e., Gaussian and Shot noise) before and after applying DiCoTTA, demonstrating that DiCoTTA effectively reduces the gap between different domains. In Fig. 3(b), we quantitatively measured the Chamfer distance between domain embeddings from different domains before and after applying DiCoTTA. This result indicates that the distance decreases as intended after the domain-invariant learning.
Domain embedding. In Fig. 4(a), we empirically verify that domain embeddings capture domain-specific information more prominently than content information. To assess the content information within the domain embeddings, we performed linear probing by training a linear classification layer using the domain embeddings as input while keeping the model frozen. The results show that domain classification using the domain embeddings consistently achieves higher accuracy than class classification, indicating that the domain embeddings predominantly encode domain-specific information with minimal content information, as intended.
Domain prototype selection. We qualitatively investigate how well domain prototypes represent the entire set of domain embeddings. Fig. 4(b) provides a $t$-SNE visualization of domain embeddings and selected domain prototypes, specifically conducted on the Gaussian noise domain. This visualization demonstrates that our domain prototypes are selected to be diverse and well represent the overall distribution of the domain embeddings.
Domain prototype update. Fig. 4(c) demonstrates the effectiveness of the prototype updating strategy. By forwarding the corresponding images of each prototype through the updated model, we obtain target prototypes that represent the desired prototype representations in the updated feature space. The MMD value between these target prototypes and our updated prototypes is lower than those between target prototypes and original prototypes, where MMD values are averaged across all CTTA steps. This result validates that our prototype updating method effectively moves the prototypes toward their targeted representations.
@@ 3.5 Empirical justification
Yang et al. [75] 24.0 +11.4
DiCoTTA 23.3 +12.1
Table 2: Classification error rates (%) on CIFAR100-to-CIFAR100-C online CTTA task.
@@
Yang et al. [75] 42.7 +13.1
DiCoTTA 42.5 +13.3
Table 3: Classification error rates (%) on ImageNet-to-ImageNet-C online CTTA.
@@
DiCoTTA 2.5 12.6 12.8 18.7 12.6 11.8 +13.5
Source [13] ViDA [44] Continual-MAE [43] DiCoTTA
@@
Backbone Metric Source [13] TENT [36] CoTTA [71] ViDA [44] DiCoTTA
@@
Gain / +3.3 -4.3 +4.8 +11.5
@@
Gain / +1.8 -0.1 +5.2 +22.0
Table 6: Average classification error rates (%) on CIFAR10-to-CIFAR10C CTTA task using pre-trained encoders from foundation models (DINOv2 and SAM).
[…]
Table 7: Segmentation results on Cityscape-to-ACDC online CTTA task. Mean is the average score of mIoU.
4 Experiments
4.1 Experimental setting
Evaluation benchmarks. Our method was evaluated on both classification and semantic segmentation tasks. For classification, we utilized three benchmarks [25]: CIFAR10-C, CIFAR100-C, and ImageNet-C, each consisting of 15 corruption types and five severity levels. We reported the results for the most severe level (level 5). For segmentation, we used the Cityscapes dataset [11] as the source dataset and the ACDC dataset [60] as the test dataset, including four unseen test domains: Fog, Night, Rain, and Snow.
Evaluation scenario. Following Wang et al. [71], DiCoTTA was evaluated under an online CTTA setting where the prediction is assessed immediately after data are streamed, and the pre-trained model is adapted on the fly. For the segmentation task, the test domains are repeated cyclically for three rounds. Additionally, we evaluated the generalization ability to unseen domains on CIFAR10-C using the leave-one-domain-out rule [80, 39, 44], applying CTTA methods to 10 domains of ImageNet-C and testing on the 5 unseen domains without any model adaptation. In line with recent CTTA studies [44, 43, 75], we reported results for all methods using a common ViT backbone. This follows the standardization introduced by ViDA [44], which established the use of ViTs in CTTA and re-implemented previous CNN-based methods within this ViT-based framework. For additional experiments beyond those reported by previous studies, we evaluated methods with publicly available code.
Implementation details. We utilized ViT-base [13] for classification tasks and Segformer-B5 [73] for the segmentation task as the backbone. Input images were resized to $224\times 224$ for ViT-base and to $960\times 540$ for Segformer-B5. The size of the queue storing the domain embeddings and domain prototypes was set to 256 and 40, respectively. The domain information extractor $\psi_{\textrm{domain}}$ and domain discriminator $D$ were both implemented as MLP whose activation functions are ReLU [1] and GELU [26], respectively. The domain amplifier was implemented as an adapter that includes an up-projection layer and a down-projection layer, with an intermediate dimension of 128. More details are given in Sec. B.1 of the supplement.
4.2 Evaluation on classification tasks
To evaluate the performance of DiCoTTA, we conducted extensive experiments on three conventional benchmarks: CIFAR10-C, CIFAR100-C, and ImageNet-C.
Performance on corruption benchmarks. Table 1 and Table a3 present the results on CIFAR10-C and CIFAR100-C, respectively. DiCoTTA significantly outperformed all other methods across various corruption types. Specifically, it achieved the lowest mean error rate of 7.7% on CIFAR10-C and 23.3% on CIFAR100-C, showing improvements of 20.5%p and 12.1%p over the source model, respectively. As shown in Table a6, DiCoTTA also achieved the best performance on the ImageNet-C dataset.
Generalization performance on unseen domains. Table 5 summarizes the classification error rates on 5 unseen domains of CIFAR10-C after CTTA over 10 domains. DiCoTTA achieved the lowest mean error rate, highlighting the benefit of domain-invariant learning and demonstrating its generalization ability by maintaining robustness even on unseen domains. This indicates that DiCoTTA not only adapts well to continually changing domains but also generalizes effectively to unseen domains.
Gradually changing results. Table 5 shows results of continual test-time adaptation on a CIFAR10-C gradually changing setup [71], where the corruption types changed gradually over time. DiCoTTA still demonstrated superior performance, notably outperforming existing methods.
Continual adapting for foundation models. Foundation models trained on large-scale datasets possess powerful generalization capabilities, but adapting them to downstream datasets can further improve their performance on target tasks [81, 45, 56]. To evaluate the effectiveness of our method on foundation models, we applied DiCoTTA to DINOv2 [54], a representative image-level foundation model, and SAM [33], a pixel-level foundation model, under continually changing test domains. Following the experimental setting of ViDA [44], we treated these foundation models as pre-trained models and applied each CTTA method to them on CIFAR10-C. Table 6 shows the results, demonstrating that our method enhanced the performance of both models and enabled continual adaptation without degradation.
4.3 Evaluation on segmentation tasks
Table 7 presents the segmentation results from adapting Cityscapes to ACDC, a real-world adverse condition dataset. Across three rounds, DiCoTTA consistently achieved the highest mIoU among existing CTTA methods. In the first round, our method reached an average mIoU of 62.0%, representing a 5.3%p gain over the source model. In subsequent rounds, DiCoTTA performed strongly on newly encountered domains while preserving robustness on previously encountered domains, effectively handling continually changing environments. These results suggest that DiCoTTA successfully mitigates domain shifts in semantic segmentation tasks, maintaining robustness across diverse and real-world adverse conditions. Further analyses are given in Sec. E.1 and F.3.
Error rate (%)
(a) w/o $\psi_{\textrm{amp}}$ 21.9
(b) w/ $\psi_{\textrm{amp}}$ 7.7
Table 8: Ablation study on $\psi_{\textrm{amp}}$.
Figure 5: Impact of $\psi_{\textrm{amp}}$.
@@ 4.3 Evaluation on segmentation tasks
Table 9: Ablation study on each loss of DiCoTTA.
@@ 4.4 In-depth analysis
Impact of domain amplifier. We investigated the impact of the domain amplifier by evaluating DiCoTTA with and without the domain amplifier by comparing their error rates (Table 8) and attention maps derived from the final Transformer block (Fig. 5). Removing the amplifier substantially increased error rates and produced less reliable attention maps. The results show that the domain amplifier plays a crucial role in DiCoTTA, extracting domain-specific information from each layer of the encoder, thereby enabling the encoder to become domain-invariant.
Ablation study of each loss. We conducted an ablation study to evaluate the contributions of each loss component in our method. The results, presented in Table 9, demonstrate that each loss contributes to the overall performance improvement. Notably, the domain-invariant learning loss $\mathcal{L}_{\textrm{inv}}$ was crucial for achieving high performance by guiding the model to learn domain-invariant features, which are essential for effective adaptation across different domains. Additionally, the domain discrimination loss $\mathcal{L}_{\textrm{dis}}$, which facilitated the learning of domain embeddings, was also crucial for DiCoTTA. Furthermore, both the score function for domain prototype selection $J(\mathcal{P}_{\textrm{pre}})$ and the prototype updating loss $\mathcal{L}_{\textrm{update}}$ contributed to the performance of DiCoTTA.
Sensitivity to the threshold for domain change detection. To investigate the effect of the threshold used for domain change detection, we conducted experiments varying the threshold values. Fig. 6(a) summarizes the results for thresholds ranging from 0.02 to 0.20. The performance remained stable across this range of thresholds, consistently outperforming the source model. These results suggest that DiCoTTA is insensitive to threshold settings.
Sensitivity to the queue length $|\mathcal{F}_{\text{pre}}|$. We evaluated the impact of $|\mathcal{F}_{\text{pre}}|$, which specifies the length of the queue that stores previous domain embeddings. It is designed to retain a sufficient number of embeddings while minimizing the inclusion of excessively outdated information. Fig. 6(b) presents the error rates on CIFAR10-C for six different queue lengths. The results indicate that the error rates remain consistent across all tested values of $|\mathcal{F}_{\text{pre}}|$, demonstrating that DiCoTTA is insensitive to the length of the queue. An extended analysis is provided in Sec. C.3.
Comparison of computational cost. We evaluated the computational cost of DiCoTTA in terms of inference latency and memory consumption with the prediction accuracy. As shown in Fig. 7, while TENT [68] exhibited the highest computational efficiency, it had the lowest prediction accuracy among the evaluated methods. In contrast, DiCoTTA achieved reduced inference latency and comparable memory consumption compared to other CTTA methods (excluding TENT) while outperforming them in prediction accuracy.

### v2

@@ 3.3 Domain-invariant learning
TestDG updates the model through domain-invariant learning and self-training. For domain-invariant learning, we employ a two-step process. In the first step, the domain information extractor and amplifier are updated to extract domain-specific information, i.e., domain embeddings, from the encoder features. In the second step, the encoder is updated to be domain-invariant by aligning the current domain embeddings with the previous domain prototypes (Sec. 3.2). Alternating these two steps allows the encoder to gradually become domain-invariant, and the domain information extractor and amplifier to become sensitive to the remaining domain-specific information of the encoder. In addition, following previous work [67, 17, 43], we update our model by a self-training loss $\mathcal{L}_{\textrm{self}}$, a cross-entropy loss between our model’s prediction $\hat{y}$ and a pseudo label $\tilde{y}$ generated by its exponential moving average [62]:
@@
where $C$ is the total number of classes. For inference, we use $\hat{y}$ as the classification prediction. Details of the two steps are described below.
Step 1: Extracting domain embeddings. To maintain the semantic information of the encoder features during domain-invariant learning, we first extract their domain-specific information in the form of embedding vectors. To this end, we introduce the domain information extractor that extracts domain embeddings from output features of the encoder, and the domain amplifier that amplifies the domain-specific information of the encoder to facilitate the extractor. Let $\mathcal{F}_{\textrm{cur}}=\{f_{i}\}_{i=1}^{n}$ be the domain embeddings of test input images from the current test domain and $\mathcal{P}_{\textrm{pre}}$ the previous domain’s prototypes. The domain information extractor and amplifier learn a domain embedding space by distinguishing $\mathcal{F}_{\textrm{cur}}$ and $\mathcal{P}_{\textrm{pre}}$ with a domain discrimination loss:
$\mathcal{L}_{\textrm{dis}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})=-\frac{1}{n}\bigg{\{}\displaystyle\sum_{f_{i}\in\mathcal{F}_{\textrm{cur}},}$ $\,\log(D(f_{i}))-\sum_{p_{i}\in\mathcal{P}_{\textrm{pre}}}\log(1-D(p_{i}))\bigg{\}},$ (5)
@@
Figure 3: Analysis on the impact of domain-invariant learning in TestDG. (a) 2D visualization of the distribution of domain embeddings from different domains before and after TestDG. (b) The domain gap between different domains measured by Chamfer distance before and after applying TestDG.
Figure 4: Analysis on domain embeddings and prototypes. (a) Analysis of class classification and domain classification accuracy using domain embeddings during CTTA iterations. (b) 2D visualization of domain embeddings and domain prototypes. (c) MMD between the target and original prototypes before the update and that between the target and updated prototypes.
Step 2: Learning domain-invariant features. The second step aims to optimize the encoder to achieve domain invariance across diverse test domains alongside classification (Eq. 4). TestDG accomplishes this by aligning the current domain embeddings and the previous domain’s prototypes. To be specific, we update the encoder and its classification head while freezing the domain information extractor and amplifier. Given current domain embeddings $\mathcal{F}_{\textrm{cur}}$ and previous domain prototypes $\mathcal{P}_{\textrm{pre}}$, the domain-invariant learning loss is given by
@@
By minimizing this loss, the encoder’s features are trained to produce domain embeddings that are indistinguishable between different domains, making them domain-generalizable to future domains.
@@ 3.4 Domain prototype update
Since the model used to extract the domain prototypes is constantly updated during CTTA, the domain prototypes extracted from the earlier model become outdated. Thus, the domain prototypes should also be updated to reflect the update of the model, approximating prototypes extracted from the updated model. Inspired by Asadi et al. [2], we aim to preserve the set similarity between the previous domain prototypes $\mathcal{P}_{\textrm{pre}}$ and the current domain embeddings $\mathcal{F}_{\textrm{cur}}$ to effectively keep the relationship between them intact. In specific, we measure the Chamfer distance between the two sets $\mathcal{P}_{\textrm{pre}}$ and $\mathcal{F}_{\textrm{cur}}$; a detailed analysis is provided in the supplement. For iteration $t$, we update the domain prototypes $\mathcal{P}_{\textrm{pre}}^{t-1}$ to $\mathcal{P}_{\textrm{pre}}^{t}$ to ensure that the Chamfer distance between previous domain prototypes and current domain embeddings remains consistent before and after the model update. Specifically, the domain prototypes $\mathcal{P}_{\textrm{pre}}^{t}$ are updated by minimizing the prototype updating loss:
@@
$d_{\textrm{CD}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})=\sum_{f_{i}\in\mathcal{F}_{\textrm{cur}}}\min_{p_{j}\in\mathcal{P}_{\textrm{pre}}}\left\|p_{j}-f_{i}\right\|_{2}^{2}+\sum_{p_{j}\in\mathcal{P}_{\textrm{pre}}}\min_{f_{i}\in\mathcal{F}_{\textrm{cur}}}\left\|p_{j}-f_{i}\right\|_{2}^{2}.$ (8)
3.5 Empirical justification
Domain-invariant learning. We analyze the impact of domain-invariant learning both qualitatively and quantitatively. Figure 3(a) provides a $t$-SNE visualization [63] of domain embeddings from different domains (i.e., Gaussian and Shot noise) before and after applying TestDG, demonstrating that TestDG effectively reduces the gap between different domains. Figure 3(b) shows the Chamfer distance between domain embeddings from different domains before and after applying TestDG, indicating that the distance decreases as intended after the domain-invariant learning.
Domain embedding. In Figure 4(a), we empirically verify that domain embeddings capture domain-specific information more prominently than content information. To assess the content information within the domain embeddings, we performed linear probing by training a linear classification layer using the domain embeddings as input while keeping the model frozen. The results show that domain classification using the domain embeddings consistently achieves higher accuracy than class classification, indicating that the domain embeddings predominantly encode domain-specific information with minimal content information, as intended.
Domain prototype selection. We qualitatively investigate how well domain prototypes represent the entire set of domain embeddings. Figure 4(b) provides a $t$-SNE visualization of domain embeddings and selected domain prototypes on the Gaussian noise domain, showing that our domain prototypes are diverse and well represent the overall distribution of the domain embeddings.
Domain prototype update. Figure 4(c) demonstrates the effectiveness of the prototype updating strategy. By forwarding the corresponding images of each prototype through the updated model, we obtain ideal target prototypes in the updated feature space. The MMD value between these target prototypes and our updated prototypes is lower than those between target prototypes and original prototypes, where MMD values are averaged across all CTTA steps. This result validates that our updating method effectively moves the prototypes toward their targeted representations.
4 Experiments
4.1 Experimental setting
Evaluation benchmarks. Our method was evaluated on both classification and semantic segmentation tasks. For classification, we utilized three benchmarks [25]: CIFAR10-C, CIFAR100-C, and ImageNet-C, each consisting of 15 corruption types and five severity levels. We reported the results for the most severe level (level 5). For segmentation, we used Cityscapes [11] as the source and ACDC [56] as a test dataset, including four unseen test domains: Fog, Night, Rain, and Snow.
Evaluation scenario. Following [67], TestDG was evaluated under an online CTTA setting where the prediction is assessed immediately after data are streamed, and the pre-trained model is adapted on the fly. For the segmentation task, test domains are repeated cyclically for three rounds. Additionally, we evaluated the generalization ability to unseen domains on CIFAR10-C using the leave-one-domain-out rule [76, 38, 43], applying CTTA methods to 10 domains of ImageNet-C and testing on the 5 unseen domains without any model adaptation. In line with recent CTTA studies [43, 42, 71], we reported results for all methods using a common ViT backbone. This follows the standardization introduced by ViDA [43], which established the use of ViTs in CTTA and re-implemented previous CNN-based methods within this ViT-based framework. For additional experiments beyond those reported by previous studies, we evaluated methods with publicly available code.
Implementation details. We utilized ViT-base [13] for classification tasks and Segformer-B5 [69] for the segmentation task as the backbone. Input images were resized to $384\times 384$ for CIFAR10-C and CIFAR100-C, $224\times 224$ for ImageNet-C, and to $960\times 540$ for ACDC. The size of the queue storing the domain embeddings and domain prototypes was set to 256 and 40, respectively. The domain information extractor $\psi_{\textrm{domain}}$ and domain discriminator $D$ were both implemented as MLP whose activation functions are ReLU [1] and GELU [26], respectively. The domain amplifier was implemented as an adapter that includes an up-projection layer and a down-projection layer, with an intermediate dimension of 128. More details are given in Sec. B.1 of the supplement.
@@ 4.1 Experimental setting
Table 2: Classification error rates (%) on CIFAR100-to-CIFAR100-C CTTA.
@@
Versatile [71] 24.0 +11.4
TestDG 23.3 +12.1
Table 3: Classification error rates (%) on ImageNet-to-ImageNet-C CTTA.
@@
Versatile [71] 42.7 +13.1
TestDG 42.5 +13.3
@@
TestDG 2.5 12.6 12.8 18.7 12.6 11.8 +13.5
Source [13] ViDA [43] Continual-MAE [42] TestDG
@@
Table 6: Segmentation mIoU results on Cityscape-to-ACDC online CTTA task.
@@
4.2 Evaluation on classification tasks
To evaluate the performance of TestDG, we conducted extensive experiments on three conventional benchmarks: CIFAR10-C, CIFAR100-C, and ImageNet-C.
Performance on corruption benchmarks. Table 1 and Table 5 present the results on CIFAR10-C and CIFAR100-C. TestDG significantly outperformed all other methods across various corruption types. Specifically, it achieved the lowest mean error rate of 7.7% on CIFAR10-C and 23.3% on CIFAR100-C, showing improvements of 20.5%p and 12.1%p over the source model, respectively. As shown in Table 5, TestDG also achieved the best performance on the ImageNet-C dataset, where performance differences are naturally smaller due to the large number of classes.
Generalization performance on unseen domains. Table 5 summarizes the classification error rates on 5 unseen domains of CIFAR10-C after CTTA over 10 domains. TestDG achieved the lowest mean error rate, highlighting the benefit of domain-invariant learning and demonstrating its generalization ability by maintaining robustness even on unseen domains. This indicates that TestDG not only adapts well to continually changing domains but also generalizes effectively to unseen domains.
Gradually changing results. Table 5 shows results of continual test-time adaptation on a CIFAR10-C gradually changing setup [67], where the corruption types changed gradually over time. TestDG still demonstrated superior performance, notably outperforming existing methods.
4.3 Evaluation on segmentation tasks
Table 6 presents the segmentation results from adapting Cityscapes to ACDC, a real-world adverse condition dataset. Across three rounds, TestDG consistently achieved the highest mIoU among existing CTTA methods. In the first round, our method reached an average mIoU of 62.0%, representing a 5.3%p gain over the source model. In subsequent rounds, TestDG performed strongly on newly encountered domains while preserving robustness on previously encountered domains, effectively handling continually changing environments. These results suggest that TestDG successfully mitigates domain shifts in semantic segmentation tasks, maintaining robustness across diverse and real-world adverse conditions. Further analyses are given in Sec. E.1 and F.3.
Table 7: Ablation study on each component of TestDG.
@@ 4.3 Evaluation on segmentation tasks
21.9
7.7
@@ 4.4 In-depth analysis
Impact of domain amplifier. We investigated the impact of the domain amplifier by evaluating TestDG with and without the domain amplifier by comparing their error rates (Table 4.3). The results show that the domain amplifier plays a crucial role in TestDG, extracting domain-specific information from each layer of the encoder, thereby enabling the encoder to become domain-invariant.
Ablation study of each loss. We conducted an ablation study to evaluate the contributions of each loss component in our method. The results, presented in Table 4.3, demonstrate that each loss contributes to the overall performance improvement. Notably, the domain-invariant learning loss $\mathcal{L}_{\textrm{inv}}$ was crucial for achieving high performance by guiding the model to learn domain-invariant features, which are essential for effective adaptation across different domains. Additionally, the domain discrimination loss $\mathcal{L}_{\textrm{dis}}$, which facilitated the learning of domain embeddings, was also crucial for TestDG. Furthermore, both the score function for domain prototype selection $J(\mathcal{P}_{\textrm{pre}})$ and the prototype updating loss $\mathcal{L}_{\textrm{update}}$ contributed to the performance of TestDG.
Sensitivity to the threshold for domain change detection. To investigate the effect of the threshold used for domain change detection, we conducted experiments varying the threshold values. Figure 5(a) summarizes the results for thresholds ranging from 0.02 to 0.20. The performance remained stable across this range of thresholds, consistently outperforming the source model. These results suggest that TestDG is insensitive to threshold settings.
Sensitivity to the queue length $|\mathcal{F}_{\text{pre}}|$. We evaluated the impact of $|\mathcal{F}_{\text{pre}}|$, which specifies the length of the queue that stores previous domain embeddings. It is designed to retain a sufficient number of embeddings while minimizing the inclusion of excessively outdated information. Figure 5(b) presents the error rates on CIFAR10-C for six different queue lengths. The results indicate that the error rates remain consistent across all tested values of $|\mathcal{F}_{\text{pre}}|$, demonstrating that TestDG is insensitive to the length of the queue. An extended analysis is provided in Sec. C.3.
Comparison of computational cost. We evaluated the computational cost of TestDG in terms of inference latency and memory consumption with the prediction accuracy. As shown in Figure 6, while TENT [64] exhibited the highest computational efficiency, it had the lowest prediction accuracy among the evaluated methods. In contrast, TestDG achieved reduced inference latency and comparable memory consumption compared to other CTTA methods (excluding TENT) while outperforming them in prediction accuracy.